Semigroups

Definition

A semigroup is a set S with an operation, which we will call “+” and which we refer to as “addition.” But don’t be fooled. it might be addition, but it might not be. Examples of semigroups. numbers (whatever kind)

addition

numbers (whatever kind) multiplication numbers (whatever kind)

max

numbers (whatever kind)

min

Boolean  and
Boolean  or
Boolean  exclusive or

Strings concatenation